OpenStudy (anonymous):

For which value(s) of m is y=x^m a solution to the DE x^2y''-y=0?

hartnn (hartnn):

u can first differentiate y, twice and find y'' what u get?

hartnn (hartnn):

then check whether x^2 y'' - y comes out to be 0, if its 0, then its the solution

OpenStudy (anonymous):

I tried dat..let me try again ...by the way am asked for the specic m values...

hartnn (hartnn):

oh, yes, u will get a quadratic in m when u equate that to 0, and u will get specific values of m

hartnn (hartnn):

did u get y" as m(m-1)x^(m-2) = m(m-1)x^m/x^2......?

OpenStudy (anonymous):

I got x^2((m-1)(m)x^(m-2))-x^m=0........After differentiation and substitution!!

OpenStudy (anonymous):

then add the exponents of x^2 and x^(m-2)

OpenStudy (anonymous):

=x^m

hartnn (hartnn):

x^(m-2) = x^m/x^2 so x^2 * x^(m-2) = x^m

OpenStudy (anonymous):

then take x^m common

hartnn (hartnn):

so finally did u get m^2-m-1=0 can u solve this ?

OpenStudy (anonymous):

=x^m x^(m-2) = x^m/x^2 so x^2 * x^(m-2) = x^m

OpenStudy (anonymous):

i think that is true answer

OpenStudy (anonymous):

but u cannot solve for m to make the equation =0 canu?

hartnn (hartnn):

u can solve his m^2-m-1=0 or u have some other doubt ?

OpenStudy (anonymous):

cool...so it's not a whole number I see...

hartnn (hartnn):

yup, its not a whole number.

OpenStudy (anonymous):

u want a medal?lol

hartnn (hartnn):

if u get the right answer.....

OpenStudy (anonymous):

:)

hartnn (hartnn):

thank u :)